In nine years, Chad will be three times my present age. I will then be half as old as he. How old am I now?
Please show steps on how to solve this
Thanks

Answer by stanbon(75887)

In nine years, Chad will be three times my present age. I will then be half as old as he. How old am I now?
-------
Equations:
c+9 = 3m
m+9 = (1/2)(c+9)
-------
Notice:: In the 2nd equation, (c+9) = 2(m+9)
----
Equation:
3m = 2(m+9)
m = 18 (My age now)
---
Cheers,
Stan H.
----------